Suppose z0 is any constant complex number interior to any simple closed

contour C. Show that for a positive integer n,

dz

(z−z0)

n = {

2πi, n = 1

C 0, n > 1.


1
Expert's answer
2022-01-03T17:18:02-0500

Cauchy’s integral formula for derivatives:

"f^{(n)}(z_0)=\frac{n!}{2\pi i}\int \frac{f(z)}{(z-z_0)^{n+1}}dz"


we have:

"f(z)=1"


then:

for n = 1:

"f(z_0)=\frac{1}{2\pi i}\int \frac{1}{z-z_0}dz=1"


"\int \frac{1}{(z-z_0)^{n}}dz=2\pi i"


for n > 1:

"f^{(n-1)}(z)=0"


"f^{(n-1)}(z_0)=\frac{n!}{2\pi i}\int \frac{1}{(z-z_0)^{n}}dz=0"


"\int \frac{1}{(z-z_0)^{n}}dz=0"